# Ambient space and integration of the trace anomaly

### v1

@@ Abstract
We use the ambient space construction, in which spacetime is mapped into a special lightcone of a higher dimensional manifold, to derive the integrable terms of the trace anomaly in even dimensions. We argue that the natural topological anomaly is the so-called $Q$-curvature, which, when projected from the ambient space, always comes with a Weyl covariant operator that should be adopted for the integration of the anomaly itself in the form of a nonlocal action. The use of the ambient space makes trasparent the fact that there are some new ambiguities in the integration of the anomaly, which we now understand geometrically from the ambient point of view. These ambiguities, which manifest themselves as undetermined parameters in the integrated nonlocal action, become more severe in dimensions $d\geq 6$ and do not seem to be related to a choice of the renormalization scheme.
@@ I Introduction
In a path-integral approach, the quantization procedure relates the expectation value of the energy-momentum tensor $\langle T_{\mu\nu}\rangle$ to the variation of the effective action $\Gamma$ with respect to the metric, which is thus its source. The procedure introduces an anomalous nonzero value to the expectation value of the trace of the energy-momentum tensor, $\langle T\rangle\equiv g^{\mu\nu}\langle T_{\mu\nu}\rangle\neq 0$, known as trace-anomaly, assuming that diffeomorphisms invariance (and thus covariance) is preserved during quantization Duff:1993wm . The trace anomaly cannot be removed by changing the regularization scheme of the path-integral and, for this reason, it is a special property of the CFT that we have to live with. In fact, the coefficients of the anomaly manifest themselves through correlators that involve multiple copies of the energy-momentum tensors, as well as the other relevant operators of the theory, also in the flat space limit.
@@ II.2 Ambient space, obstructions and holography
A consequence of the Ricci-flatness condition is that ${\rm Ric}_{d+1}\sim g_{d+1}$ on the boundary, so, asymptotically, they approach AdS spaces. The ambient space is thus referred to as an asymptotically local Anti-de Sitter space (ALAdS) Parisini:2022wkb ; Parisini:2023nbd .
@@ II.6 PBH diffeomorphisms as generators of conformal transformations
III The integrable terms for the trace-anomaly
@@
Before diving into the computations, it is useful to quickly recapitulate through one comprehensive example what we are looking for. We choose the $d=4$ trace anomaly because it brings forth all the important ingredients that appear in the general case. Consider a classical Weyl-invariant action $S[\varphi]$ of a Weyl-covariant field $\varphi$ coupled to a nondynamical metric. The Noether identities tell us that on-shell the trace of the variational energy-momentum tensor, $T^{\mu\nu}=-\frac{2}{\sqrt{g}}\frac{\delta S}{\delta g_{\mu\nu}}$, is zero, i.e., $T^{\mu}{}_{\mu}=0$. Putting the action in a path-integral and integrating in $\varphi$ defines the effective action $\Gamma$, from which one can compute the expectation value $\langle T^{\mu\nu}\rangle=-\frac{2}{\sqrt{g}}\frac{\delta\Gamma}{\delta g_{\mu\nu}}$, being $g_{\mu\nu}$ the source of the operator $T^{\mu\nu}$ (e.g., see Ref. Hathrell:1981zb for a self-interacting scalar). The path-integral induces an anomaly of Weyl invariance, known as trace anomaly Duff:1977ay , which, at fixed points of the renormalization group, can be computed using the method outlined in Appendix C and is of the form
@@
The anomaly is constrained by a Wess-Zumino consistency condition $[\delta_{\sigma},\delta_{\sigma^{\prime}}]\Gamma=0$ Osborn:1991gm , which can be structured as a cohomological problem Bonora:1983ff ; Bonora:1985cq . The consistency is reflected in integrability conditions for the anomaly’s constituents, that is, $[\delta_{\sigma},\delta_{\sigma^{\prime}}]E_{4}=0$ and $[\delta_{\sigma},\delta_{\sigma^{\prime}}]W^{2}=0$. The structure of $\langle T\rangle$ is general in that we expect that in any even dimension $d=2n$ the anomaly is the sum of three contributions, a topological $a$-anomaly, a Weyl covariant $b$-anomaly Cardy:1988cwa ; Myers:2010xs , and a trivial $a^{\prime}$-anomaly which is a total derivative. All these charges have and have been regarded as potential generalizations of the central charge $c$ of Zamolodchikov’s theorem Zamolodchikov:1986gt , for example see Ref. Duff:1993wm for context and Refs. Deser:1993yx ; Cappelli:1990yc ; Shore:1990wq ; Osborn:1991gm ; Anselmi:1999xk ; Anselmi:1999uk for examples, though the topological anomaly is generally regarded as the best candidate for irreversibility.
@@ IV Integration of the anomaly in any $d$
In this section we choose $d$ even, i.e., $d=2n$. Consequently, all the tensors coming from the ambient space construction are well-defined up to a certain order in $p$ of the expansion (18). For the purpose of this section, it is sufficient that the objects $Q_{d}$ and $\Delta_{d}$, introduced in Sect. III.1 as parts of the differential operators $D_{2n,d}$, and $W_{d,I}$, introduced in Sect. III.2, are well-defined, which is in fact true and has been checked by explicit computations in $d=2,4,6$. As hinted at the beginning of Sect. III, the general form of the trace-anomaly in $d=2n$ is expected to be the sum of three terms
@@
which are referred to as $b$-anomalies, $a$-anomaly and trivial (or $a^{\prime}$-) anomalies, respectively Deser:1993yx ; Duff:1993wm . The terms of the anomaly come from the construction of a path-integral and depend on the conformal fields that are integrated-out. A relatively general procedure for obtaining $\langle T\rangle$ is sketched in Appendix C, in which we also show a basic set of assumptions that lead to the desired result.
@@
$E_{d}(g)=\frac{1}{2^{n}}R_{\mu_{1}\mu_{2}\nu_{1}\nu_{2}}\cdots R_{\mu_{n-1}\mu_{n}\nu_{n-1}\nu_{n}}\epsilon^{\mu_{1}\cdots\mu_{n}}\epsilon^{\nu_{1}\cdots\nu_{n}}\,.$ (98)
@@
it is sufficient to change the renormalization scheme of the effective action subtracting $\Gamma\to\Gamma^{\prime}=\Gamma-F$ to obtain a new energy-momentum tensor that does not include the last term. This is of course possible only in a specific scheme and because the metric is not dynamical. Notice that it is straightforward to prove that the variation of a scale-but-not-conformally invariant functional $F$ must be the total derivative of a certain vector Nakayama:2016dby , while the opposite may not be obvious, although always true to our knowledge. Nevertheless, in general there are actually multiple options for the functional $F[g]$ given any $J^{\mu}_{d}$, which is another manifestation of the scheme dependence.
@@
In general $\Gamma_{\rm an,loc}$ depends on the chosen scheme, inherited by the form of $J^{\mu}_{d}$.
@@
where the first term contains any possible local/nonlocal conformal action (undetermined by the anomaly), the second term is a scheme-dependent contribution, and the third nonlocal term is entirely constrained by the form of the trace-anomaly and generalizes the Polyakov’s and Riegert’s actions Polyakov:1981rd ; Riegert:1984kt . The coefficients $\tilde{a}$ and $\tilde{b}_{I}$ can be given in terms of the original $a$ and $b_{I}$ as outlined for Eq. (103). The nonlocal part of the above action can be “localized” introducing at least two auxiliary fields Balbinot:1999ri , however, if we exploit the fact that the conformally invariant part is undetermined, we can actually “complete” the square and write it as
@@ V.1 Nonlocal action in $d=6$ and ambiguities
$\langle T\rangle_{a^{\prime}}\equiv\nabla_{\mu}J^{\mu}_{6}=\sum_{i=1}^{7}f_{i}B^{i}\,,$ (141)
where the $B^{i}$ are seven independent boundary terms of dimension six that can be constructed, as detailed in Appendix B. The topological $a$-anomaly is given by the six-dimensional Euler density $E_{6}$. Combining everything, we can express the vacuum expectation value of the trace of the energy-momentum tensor in the standard form as
@@
$J^{\alpha}_{6}=$ $-\left(\frac{408}{5}-20\zeta\right)R^{\nu}{}_{\mu}\nabla_{\nu}R^{\nu\alpha}-\left(\frac{36}{25}-2\zeta\right)R^{\alpha\mu}\nabla_{\mu}R+\zeta\nabla^{\alpha}(R^{2})$ (144)
@@
$\langle T\rangle=a_{1}\tilde{E}_{6}+\sum_{i=1}^{7}{f^{\prime}_{1}}_{i}B^{i}\,,$ (145)
where the $B_{i}$ are the same total derivative terms, as mentioned earlier. As discussed also in Anselmi:1999uk , this expression can be readily integrated on conformally flat spacetimes, where $\sqrt{g}\langle T\rangle\propto\partial^{6}\sigma$, by exploiting the conformal properties of the pondered Euler density. Through straightforward manipulations, one can show that a relation similar to Eq. (143) holds also between $Q_{6}$ and $\tilde{E}_{6}$. Moreover, from their definitions, it is simple to see that both quantities transform identically under conformal transformations when the metric is decomposed as $g_{\mu\nu}=e^{2\sigma}\delta_{\mu\nu}$.
@@
$E_{6}-6\,Q_{6}+\sum_{i=1}^{7}f_{i}B_{i}+\sum_{I=1}^{3}c_{I}W_{6,I}=0\,,$ (146)
which holds for the following constants
@@
$\langle T\rangle=\sum_{I=1}^{3}\tilde{b}_{I}W_{6,I}+\tilde{a}Q_{6}+\sum_{i=1}^{7}\tilde{f}_{i}B^{i}\,,$ (148)
where $\tilde{a}=6a$, and the constants $\tilde{b}_{I}$ and $\tilde{f}_{i}$ are obtained by constant shifts according to Eqs. (147). Once again, using the known covariance properties of $Q_{6}$, the anomaly becomes easily integrable in this form.
@@
Thus the anomaly can also be expressed as $\langle T\rangle=\sum_{I=1}^{3}\tilde{b}_{I}W_{6,I}+\tilde{a}\mathcal{P}_{6}+\sum_{i=1}^{7}\tilde{f}_{i}B^{i}$, but the charges now differ from the previous choice in that they depend on the parameters $\alpha_{i}$. Following the usual steps, we find the nonlocal part of the anomalous action
@@ Acknowledgments.
We are grateful to Davide Pittet for collaboration on topics related to this project during the development of his MSc thesis on ambient space and Ricci-gauging at the University of Pisa pittet . OZ is also grateful to Andreas Stergiou for early clarifications on the role of $Q$-curvatures that instigated the beginning of this work. We are also grateful to Claudio Corianò for important discussions that motivated us to conclude this project.

### v2

@@ Abstract
We use the ambient space construction, in which spacetime is mapped into a special lightcone of a higher dimensional manifold, to derive the integrable terms of the trace anomaly in even dimensions. We argue that the natural topological anomaly is the so-called $Q$-curvature, which, when projected from the ambient space, always comes with a Weyl covariant operator that can naturally be adopted for the integration of the anomaly itself in the form of a nonlocal action. The use of the ambient space makes trasparent the fact that there are some new ambiguities in the integration of the anomaly, which we now understand geometrically from the ambient point of view. These ambiguities, which manifest themselves as undetermined parameters in the integrated nonlocal action, become more severe in dimensions $d\geq 6$ and do not seem to be related to a choice of the renormalization scheme.
@@ I Introduction
In a path-integral approach, the quantization procedure relates the expectation value of the energy-momentum tensor $\langle T_{\mu\nu}\rangle$ to the variation of the effective action $\Gamma$ with respect to the metric, which is thus its source. The procedure introduces an anomalous nonzero value to the expectation value of the trace of the energy-momentum tensor, $\langle T\rangle\equiv g^{\mu\nu}\langle T_{\mu\nu}\rangle\neq 0$, known as trace anomaly, assuming that diffeomorphisms invariance (and thus covariance) is preserved during quantization Duff:1993wm . The trace anomaly cannot be removed by changing the regularization scheme of the path-integral and, for this reason, it is a special property of the CFT that we have to live with. In fact, the coefficients of the anomaly manifest themselves through correlators that involve multiple copies of the energy-momentum tensors, as well as the other relevant operators of the theory, also in the flat space limit.
@@ II.2 Ambient space, obstructions and holography
A consequence of the Ricci-flatness condition is that ${\rm Ric}_{d+1}\sim g_{d+1}$ on the boundary, so, asymptotically, they approach AdS spaces. The ambient space is thus referred to as an asymptotically local Anti-de Sitter space (ALAdS) Parisini:2022wkb ; Parisini:2023nbd . Given the intimate connection with holography, the formalism that we are about to flesh out can then be used to compute the holographic anomaly (see Refs. Bugini:2018def ; Jia:2021hgy and references within).
@@ II.6 PBH diffeomorphisms as generators of conformal transformations
III The integrable terms for the trace anomaly
@@
Before diving into the computations, it is useful to quickly recapitulate through one comprehensive example what we are looking for. We choose the $d=4$ trace anomaly because it brings forth all the important ingredients that appear in the general case. Consider a classical Weyl-invariant action $S[\varphi]$ of a Weyl-covariant field $\varphi$ coupled to a nondynamical metric. The Noether identities tell us that on-shell the trace of the variational energy-momentum tensor, $T^{\mu\nu}=-\frac{2}{\sqrt{g}}\frac{\delta S}{\delta g_{\mu\nu}}$, is zero, i.e., $T^{\mu}{}_{\mu}=0$. Putting the action in a path-integral and integrating in $\varphi$ defines the effective action $\Gamma$, from which one can compute the expectation value $\langle T^{\mu\nu}\rangle=-\frac{2}{\sqrt{g}}\frac{\delta\Gamma}{\delta g_{\mu\nu}}$, being $g_{\mu\nu}$ the source of the operator $T^{\mu\nu}$ (e.g., see Ref. Hathrell:1981zb for a self-interacting scalar). The path-integral induces an anomaly of Weyl invariance, known as trace anomaly Duff:1977ay , which, at fixed points of the renormalization group, can be computed using the method outlined in Appendix D and is of the form
@@
The anomaly is constrained by a Wess-Zumino consistency condition $[\delta_{\sigma},\delta_{\sigma^{\prime}}]\Gamma=0$ Osborn:1991gm , which can be structured as a cohomological problem Bonora:1983ff ; Bonora:1985cq or even more Jia:2023tki . The consistency is reflected in integrability conditions for the anomaly’s constituents, that is, $[\delta_{\sigma},\delta_{\sigma^{\prime}}]E_{4}=0$ and $[\delta_{\sigma},\delta_{\sigma^{\prime}}]W^{2}=0$. The structure of $\langle T\rangle$ is general in that we expect that in any even dimension $d=2n$ the anomaly is the sum of three contributions, a topological $a$-anomaly, a Weyl covariant $b$-anomaly Cardy:1988cwa ; Myers:2010xs , and a trivial $a^{\prime}$-anomaly which is a total derivative constrained by the Wess-Zumino consistency condition. [^8] In $d=4$ the only trivial anomaly is $\Box_{g}R$ and satisfies the integrability condition, but in $d=6$ and higher total derivatives are not guaranteed to always come from Weyl variations of local terms Bastianelli:2000rs . All these charges have and have been regarded as potential generalizations of the central charge $c$ of Zamolodchikov’s theorem Zamolodchikov:1986gt , for example see Ref. Duff:1993wm for context and Refs. Deser:1993yx ; Cappelli:1990yc ; Shore:1990wq ; Osborn:1991gm ; Anselmi:1999xk ; Anselmi:1999uk for examples, though the topological anomaly is generally regarded as the “best” candidate for irreversibility.
@@ III.1 Powers of the ambient Laplacian: GJMS hierarchy and $Q$-curvatures
The only subtlety in the above logic involves the total derivative, which does not manifestly satisfies the Wess-Zumino integrability condition. This issue is examined in more detail in Sect. IV. Nevertheless, we anticipate that the Weyl covariance properties of all the terms involved in the parametrization of the anomaly ensures that the total derivative term also inherently satisfies the Wess-Zumino consistency conditions.
@@ IV Integration of the anomaly in any $d$
In this section we choose $d$ even, i.e., $d=2n$. Consequently, all the tensors coming from the ambient space construction are well-defined up to a certain order in $p$ of the expansion (18). For the purpose of this section, it is sufficient that the objects $Q_{d}$ and $\Delta_{d}$, introduced in Sect. III.1 as parts of the differential operators $D_{2n,d}$, and $W_{d,I}$, introduced in Sect. III.2, are well-defined, which is in fact true and has been checked by explicit computations in $d=2,4,6$. As hinted at the beginning of Sect. III, the general form of the trace anomaly in $d=2n$ is expected to be the sum of three terms
@@
which are referred to as $b$-anomalies, $a$-anomaly and trivial (or $a^{\prime}$-) anomalies, respectively Deser:1993yx ; Duff:1993wm . The terms of the anomaly come from the construction of a path-integral and depend on the conformal fields that are integrated-out. A relatively general procedure for obtaining $\langle T\rangle$ is sketched in Appendix D, in which we also show a basic set of assumptions that lead to the desired result.
@@
$E_{d}(g)=\frac{1}{2^{n}}R_{\mu_{1}\mu_{2}\nu_{1}\nu_{2}}\cdots R_{\mu_{2n-1}\mu_{2n}\nu_{2n-1}\nu_{2n}}\epsilon^{\mu_{1}\cdots\mu_{2n}}\epsilon^{\nu_{1}\cdots\nu_{2n}}\,.$ (98)
@@
it is sufficient to change the renormalization scheme of the effective action subtracting $\Gamma\to\Gamma^{\prime}=\Gamma-F$ to obtain a new energy-momentum tensor that does not include the last term. This is of course possible only in a specific scheme and because the metric is not dynamical. More precisely, we assume that the trivial anomaly $\nabla_{\mu}J_{d}^{\mu}$ is constrained to satisfy the Wess-Zumino consistency conditions, ensuring its integrability.
Notice that it is straightforward to prove that the variation of a scale-but-not-conformally invariant functional $F$ must be the total derivative of a certain vector Nakayama:2016dby , while the opposite may not be obvious. The first nontrivial counterexample of a total derivative that is not integrable appears in $d=6$, as discussed in detailed in Ref. Bastianelli:2000rs . The analysis shows that there are seven distinct six-dimensional boundary terms can be constructed, but the space of trivial anomalies satisfying the consistency conditions is spanned by linear combinations of only six combinations. The explicit forms of these total derivatives are presented in Appendix B. That said, in general there are multiple options for the functional $F[g]$ that integrates the total derivative given any $J^{\mu}_{d}$, which is another manifestation of the scheme dependence.
@@
In general $\Gamma_{\rm an,loc}$ depends on the chosen scheme, inherited by the form of $J^{\mu}_{d}$. It is important to stress that this procedure is always applicable, as $\nabla_{\mu}\tilde{J}_{d}^{\mu}$ is constrained to satisfy the Wess-Zumino consistency conditions. As shown in Appendix C, the Euler density naturally satisfies these conditions in arbitrary even dimensions. Additionally, $Q_{d}$ and $W_{d,I}$ trivially fulfill the consistency requirements, and we have assumed from the outset that $\nabla_{\mu}{J}_{d}^{\mu}$ also complies with them (being part of the anomaly). As a result, it directly follows that $\nabla_{\mu}I_{d}^{\mu}$ and $\nabla_{\mu}\tilde{J}_{d}^{\mu}$ also satisfy the Wess-Zumino consistency conditions. To further support this point in Appendix B we explicitly verify in the first nontrivial case, which is $d=6$, that the appropriate $\nabla_{\mu}I_{d}^{\mu}$ can be derived as the variation of a local action.
@@
where the first term contains any possible local/nonlocal conformal action (undetermined by the anomaly), the second term is a scheme-dependent contribution, and the third nonlocal term is entirely constrained by the form of the trace anomaly and generalizes the Polyakov’s and Riegert’s actions Polyakov:1981rd ; Riegert:1984kt . The coefficients $\tilde{a}$ and $\tilde{b}_{I}$ can be given in terms of the original $a$ and $b_{I}$ as outlined for Eq. (103). The nonlocal part of the above action can be “localized” introducing at least two auxiliary fields Balbinot:1999ri , however, if we exploit the fact that the conformally invariant part is undetermined, we can actually “complete” the square and write it as
@@ V Generalized GJMS and integration ambiguities
While we will not consider this possibility, as it is not directly related to the trace anomaly, it is worth noting that the GJMS construction can be generalized to include fields with spin. For instance, Ref. Aros:2022ecb illustrates such a generalization by deriving higher-derivative conformal operators acting on transverse-traceless symmetric rank-2 tensors and vectors.
@@ V.1 Nonlocal action in $d=6$ and ambiguities
$\langle T\rangle_{a^{\prime}}\equiv\nabla_{\mu}J^{\mu}_{6}=\sum_{i=1}^{6}f_{i}B^{i}\,,$ (141)
where the $B^{i}$ are six independent boundary terms of dimension six that satisfy the Wess-Zumino consistency conditions, as detailed in Appendix B. The topological $a$-anomaly is given by the six-dimensional Euler density $E_{6}$. Combining everything, we can express the vacuum expectation value of the trace of the energy-momentum tensor in the standard form as
@@
$J^{\alpha}_{6}=$ $-\left(\frac{408}{5}-20\zeta\right)R^{\nu}{}_{\mu}\nabla_{\nu}R^{\mu\alpha}-\left(\frac{36}{25}-2\zeta\right)R^{\alpha\mu}\nabla_{\mu}R+\zeta\nabla^{\alpha}(R^{2})$ (144)
@@
$\langle T\rangle=a_{1}\tilde{E}_{6}+\sum_{i=1}^{6}{f^{\prime}_{1}}_{i}B^{i}\,,$ (145)
where the $B_{i}$ are the same total derivative terms mentioned earlier, evaluated in the appropriate limit. As discussed also in Anselmi:1999uk , this expression can be readily integrated on conformally flat spacetimes, where $\sqrt{g}\langle T\rangle\propto\partial^{6}\sigma$, by exploiting the conformal properties of the pondered Euler density. Through straightforward manipulations, one can show that a relation similar to Eq. (143) holds also between $Q_{6}$ and $\tilde{E}_{6}$. Moreover, from their definitions, it is simple to see that both quantities transform identically under conformal transformations when the metric is decomposed as $g_{\mu\nu}=e^{2\sigma}\delta_{\mu\nu}$.
@@
$E_{6}-6\,Q_{6}+\sum_{i=1}^{7}f_{i}A_{i}+\sum_{I=1}^{3}c_{I}W_{6,I}=0\,,$ (146)
where the $A^{i}$ are the seven possible boundary terms of dimension six reported in Appendix B. This relation holds for the following constants
@@
As previously discussed, the total derivative in the linear relation between the $Q$-curvature and the Euler density is required to satisfy the Wess-Zumino consistency conditions in each even dimension. Consequently, it must be possible to express $\sum_{i=1}^{7}f_{i}A_{i}=\sum_{i=1}^{6}k_{i}B_{i}$. We explicitly verify this fact in the $d=6$ case in Appendix B.
@@
$\langle T\rangle$ $=\sum_{I=1}^{3}\tilde{b}_{I}W_{6,I}+\tilde{a}Q_{6}+\sum_{i=1}^{6}\tilde{f}_{i}B^{i}\,,$ (148)
where $\tilde{a}=6a$, and the constants $\tilde{b}_{I}$ and $\tilde{f}_{i}$ are obtained by constant shifts according to Eqs. (147) and the results of Appendix B. Once again, using the known covariance properties of $Q_{6}$, the anomaly becomes easily integrable in this form.
@@
As before, the anomaly can be expressed as $\langle T\rangle=\sum_{I=1}^{3}\tilde{b}_{I}W_{6,I}+\tilde{a}\mathcal{P}_{6}+\sum_{i=1}^{6}\tilde{f}_{i}B^{i}$, see also Appendix B. However, the charges now differ from the previous choice in that they depend on the parameters $\alpha_{i}$. Following the usual steps, we find the nonlocal part of the anomalous action
@@ Acknowledgments.
We are grateful to Davide Pittet for collaboration on topics related to this project during the development of his MSc thesis on ambient space and Ricci-gauging at the University of Pisa pittet . OZ is also grateful to Andreas Stergiou for early clarifications on the role of $Q$-curvatures that instigated the beginning of this work. We are also grateful to Claudio Corianò for important discussions that motivated us to conclude this project and to Riccardo Martini for valuable discussions. Finally we thank Danilo Díaz Vázquez for a comment that sparked an improvement of the analysis of the trivial anomaly in the draft.